OpenStudy (anonymous):

Use the double-angle formulas to rewrite: 12cos squared x-6

OpenStudy (jdoe0001):

\(\bf 12cos^2(x-6)\quad ?\qquad 12cos^2(x)-6\quad ?\)

OpenStudy (anonymous):

\[\cos \left( a+b \right)=\cos a \cos b-\sin a \sin b\] put a=b=x \[\cos \left( x+x \right)=\cos x \cos x-\sin x \sin x\] \[\cos 2x=\cos ^{2}x-\sin ^{2}x=\cos ^{2}x-\left( 1-\cos ^{2} x\right)=2\cos ^{2}x-1\] now you can complete it.